### v1

# DiverseNet: Decision Diversified Semi-supervised Semantic Segmentation Networks for Remote Sensing Imagery

Abstract
Semi-supervised learning is designed to help reduce the cost of the manual labelling process by exploiting the utilisation of useful features from a large quantity of unlabelled data during training. Since pixel-level manual labelling in large-scale remote sensing imagery is expensive, semi-supervised learning becomes an appropriate solution to this. However, most of the existing semi-supervised learning methods still lack efficient perturbation methods to promote the precision and diversity of features (or pseudo labels) during training. In order to fill this gap, we propose DiverseNet architectures which explore multi-head and multi-model semi-supervised learning algorithms by simultaneously promoting precision and diversity during training. The two proposed methods of DiverseNet, namely the DiverseHead and DiverseModel, significantly achieve the highest semantic segmentation performance in four widely utilised remote sensing imagery data sets compared to state-of-the-art semi-supervised learning methods. Meanwhile, the proposed DiverseHead architecture is relatively lightweight in terms of parameter space compared to the state-of-the-art methods whilst reaching high-performance results for all utilised data sets.
Index Terms:
Semi-supervised Learning, Consistency Learning, Land Cover Classification
I Introduction
Supervised deep learning has gradually become the dominant technique in computer vision, especially during the last decade. Specifically, with its great success in semantic segmentation, automated land cover classification approaches have been significantly improved because of the high similarity between the two tasks [vali2020deep]. Nevertheless, supervised learning necessitates a substantial, meticulously labelled dataset. In the case of extensive remote sensing data, such as satellite imagery and drone-captured images in complex terrains, acquiring pixel-wise expert annotations proves to be a time-consuming, labour-intensive, and costly endeavour. While the field of computer vision boasts numerous well-annotated datasets, transferring deep learning models trained on these datasets to the remote sensing domain is a formidable challenge. This is primarily because of the significant disparities between natural images and remote sensing data, including hyperspectral and synthetic aperture radar (SAR) imagery, which often possesses unconventional image geometries, making the process non-intuitive. Fortunately, thanks to open-access remote sensing sources, a large amount of unlabelled imagery can be accessed easily and freely. Thus, only using a limited amount of labelled data and exploiting the benefits of a large amount of unlabelled data could be a feasible solution for solving the problem of the lack of labelled remote sensing datasets [wang2022semi].
In computer vision, semi-supervised learning has become a trending research direction to eliminate the labour-intensive and expensive annotation stages mentioned above. It leads to competitive performance compared to supervised learning in applications such as image classification [zhang2021flexmatch, huang2021embedding, wang2021semi2] and segmentation [hu2021semi, wang2022semi_u2pl, wang2021semi], which just use a part of labelled data and a large amount of unlabelled data. Specifically, mainstream semi-supervised learning methods exploit “pseudo” labels that are fake and/or auto-generated label information obtained as a result of an unsupervised step. Since pseudo labels are involved in the model training stage, the accuracy of pseudo labels is regarded as a key factor affecting the performance of semi-supervised learning approaches. However, the accuracy of pseudo labels is not always satisfactory especially when labelled training data is not complete and sufficient enough. In this case, augmenting the pseudo-labels to increase diversity has emerged as another key research direction within the semi-supervised learning domain. A greater diversity aids in constructing more comprehensive pseudo-labels, which, in turn, indirectly enhances the quality of these labels and the training effectiveness of the model.
(a) DiverseHead
(b) DiverseModel
Figure 1: Two kinds of Pseudo label generation methods based on multiple heads (a) DiverseHead and multiple models (b) DiverseModel.
In order to empower networks to harness the unexplored potentials within unlabeled data, a notable technique known as consistency regularization has emerged as a widely adopted method for semi-supervised learning. [french2019semi, filipiak2021n]. Specifically, consistency regularization methods rely on the assumption of smoothing, which posits that if two points reside in a high-density region and are close to each other, their corresponding labels should be the same or consistent [chen2010semi]. In practice, these consistency regularization methods achieve this assumption by forcing networks to produce consistent predictions for unlabeled inputs that undergo diverse perturbations.
The strategies for the above-mentioned perturbations can be categorised into three groups, namely input, feature, and network perturbations. Input perturbation is a method that involves making adjustments or alterations to the input data as part of the training process. In this way, the semi-supervised learning approach promotes the consistency of predictions made in response to these modified inputs. Widely used input perturbation methods employed in semi-supervised learning include Cutout[devries2017improved], MixUp [zhang2017mixup], and CutMix [yun2019cutmix]. Among them, CutMix stands out by combining aspects of MixUp and CutOut techniques, demonstrating superior performance compared to them [french2019semi]. However, input perturbation introduces artificial noise to the data, which might lead to incorrect or noisy labels when generating pseudo-labels for unlabeled examples. This might negatively impact the training process by providing incorrect guidance to the model when using pseudo labels. Cross-consistency training (CCT) [ouali2020semi] presents a feature perturbation method that introduces noise to both low- and high-level features. These perturbed features are then fed into multi-decoders to generate multiple outputs, followed by the enforcement of consistency among the outputs obtained from different decoders. However, feature perturbation alters the intrinsic characteristics and relationships between features, which also might lead to the creation of artificial feature representations that do not accurately reflect the underlying patterns in the original data. Consequently, the perturbed features could result in incorrect or noisy labels when generating pseudo-labels for unlabeled examples, which negatively impacts the learning process. Lastly, the network perturbation uses different networks or the same network with different initialization weights to promote diversity of predictions, such as cross pseudo supervision (CPS)[chen2021semi], confidence-guided semi-supervised learning (CGSSL)[ma2023confidence], mean teacher (MT), [tarvainen2017mean], CCT [ouali2020semi], guided collaborative training (GCT) [ke2020guided] and cross teaching between CNN and transformer [luo2022semi]. Compared to the previous two perturbation methods, although network perturbation methods introduce perturbations created by the model itself instead of artificial noise, they generally require rapidly increased computational resources due to the increased number of networks (whole or their inner stages) utilized.
I-A Contributions
The preceding conversation underscores that, although perturbation techniques have their advantages, they also bring inherent drawbacks. Therefore, the creation of methods based on perturbations is a complex task, necessitating the exploration of more sophisticated approaches to augment the potential of semi-supervised learning frameworks by incorporating advanced perturbation methods. This paper proposes perturbation-based semi-supervised network architectures, coined as DiverseNet, which consist of two novel feature and decision diversifying network architectures for the purposes of semi-supervised semantic segmentation of remote sensing imagery. First, we develop a lightweight semi-supervised semantic segmentation network called DiverseHead based on multiple decision heads to promote precision and diversity of pseudo labels whilst using two novel perturbation methods of dropout (DT) and dynamic freezing (DF). In addition, we also propose a high-performance multi-model structure called DiverseModel for scenarios equipped with high-memory computational resources. Specifically, the contributions of this work are as follows:
  1. 1.
    We promote utilisation of multiple decision heads to provide a semi-supervised learning structure, called DiverseHead. This structure is inspired by Bagging (also called bootstrap aggregating) and provides diversity for parameters and features. The main source of perturbation is from the dynamic choosing heads to be updated for each iteration. This architecture does not distinguish the main or auxiliary decoder. Also, all decoders are trained by both supervised and unsupervised losses, which comes with the advantages of high-quality perturbation and diversity of semi-supervised segmentation tasks for remote sensing imagery, whilst providing a relatively lightweight architecture.
  2. 2.
    For the purposes of creating perturbation in diversifying decisions, we ustilised two important techniques, namely ’dynamic freezing’ and ’dropout’. We make these techniques fit with DiverseHead semi-supervised learning architecture for the purpose of diversifying the parameters of the network. The two perturbation methods differ from input/feature perturbation methods, in which the source of feature diversity is from parameter diversity instead of introducing artificial noise to data or features.
  3. 3.
    A previously proposed architecture DiverseModel [ma2023confidence] has been extended for detailed evaluation and in a detailed comparison study for various semantic segmentation data sets. This architecture is based on consistency learning among multiple neural networks with distinct weights.
The rest of the paper is organised as follows: subsection I-B discusses the related work on semi-supervised semantic segmentat ion in remote sensing and some basic knowledge on ensemble machine learning whilst in Section II and III, the proposed algorithms DiverseHead and DiverseModel are presented. Section IV describes the segmentation dataset of remote sensing imagery. The experimental setting and the qualitative and quantitative analyses of the experimental results are performed in Section V. Section VI concludes the paper with a summary.
I-B Related Work
Semantic segmentation is one of the well-studied and wide-spread research areas in remote sensing society, for applications such as land cover classification/mapping [dong2019land], building change detection [zheng2022hfa], road extraction [ghandorh2022semantic], and marine debris detection [booth2023high]. Also, the number of research on semantic segmentation in remote sensing in the literature is rapidly increasing with the success of deep learning in computer vision, due to the strong task-similarity between two areas. Fully convolutional networks (FCNs) [long2015fully] is one group of the most widespread networks for segmentation tasks. They have made a considerable contribution to various segmentation tasks either in remote sensing or computer vision. Following the FCNs’ success, SegNet [badrinarayanan2017segnet] and UNet [ronneberger2015u] adopt a symmetrical encoder-decoder structure with skip connections, leveraging multi-stage features within the encoder. Alternatively, PSPNet [zhao2017pyramid] introduced a pyramid pooling structure that helps provide a global contextual understanding for pixel-level scene parsing. DeepLab architecture [chen2017deeplab] proposed a new type of convolution called atrous convolution and an atrous spatial pyramid pooling (ASPP) operation to enable the network to have the ability to control the spatial perception range of convolution kernel via setting different dilation rates. DeepLab has then been extended to DeepLabv3+ [chen2018encoder] as an improved hybrid approach that combines an encoder-decoder structure and the ASPP. Within DeepLabv3+, the ASPP module is incorporated following a ResNet backbone, enabling the network to delve into deeper layers and extract high-level features, which is helpful for accurately segmenting pixels around boundaries[xia2021cloud, zhang2019multi]. DeepLabv3+ is one of the most used networks in the literature for semi-supervised learning segmentation in the computer vision area.
Semi-supervised learning aims to alleviate the need for expensive annotation work and improve the models’ performances by making use of unlabelled data. Self-training [lee2013pseudo] (also known as pseudo labelling) represents one of the primitive semi-supervised learning strategies for both classification [kim2020distribution] and segmentation [zhu2021improving] purposes that involves the generation of pseudo labels through model predictions. These pseudo-labels are subsequently utilized to retrain the model. Various variations of self-training have been proposed focusing on determining the appropriate pseudo labels. Another widely developed semi-supervised learning approach called consistency regularization [french2019semi] is to force networks to give consistent predictions for unlabeled inputs that undergo diverse perturbations. CCT [ouali2020semi] employs an encoder-decoder architecture with multiple auxiliary decoders. These decoders introduce diversity in the output by feature perturbations specific to each auxiliary decoder. They calculate MSE loss between predictions of the main decoder and each auxiliary decoder without creating pseudo labels. It is worth noting that the unsupervised loss is not used to supervise the main decoder. Following CCT, subsequent methods like GCT [ke2020guided] and CPS [chen2021semi] are proposed to introduce network perturbation for consistency regularization. Specifically, they use two same network structures but with different weight initialization. CPS differs from GCT by employing pseudo labels to enforce consistency, whereas GCT achieves consistency regularization by utilizing predictions of networks cooperating with a flow detector. Although the network perturbation achieved through different weight initialization in the previous two works shows a good ability to generate diversity, we believe that employing different networks could yield even greater diversity.
Ensemble machine learning is a concept that employs multiple learners and combines their predictions [sewell2008ensemble]. Bagging, a form of ensemble learning, is a technique aimed at reducing prediction variance by creating multiple iterations of a predictor and then utilizing them to form an aggregated predictor [breiman1996bagging]. Specifically, Bagging creates sample subsets by randomly selecting from the training dataset and subsequently utilizes these acquired subsets to train the foundational models for integration. When forecasting a numerical result, the aggregation involves taking an average of the versions, while predicting a class, it relies on a majority vote. Based on the concept, bagging is a commonly employed approach for enhancing the robustness and precision of machine learning algorithms in terms of classification and regression applications [ren2016ensemble].
II Cross-head Supervision (DiverseHead)
Figure 2: DiverseHead: an online semi-supervised learning approach. The slide windows of the freezer randomly select a certain number of heads to freeze the parameter update of heads. A mainstream head is randomly selected to be supervised by unsupervised loss.
This section introduces the details of the proposed DiverseHead Semi-Supervised Learning method. Inspired by Bagging (also called bootstrap aggregating), DiverseHead is proposed to create various predictions by multiple heads with even the same structure. The source of perturbation is from the dynamic choosing heads to be updated for each iteration, called dynamic freezing, which is equivalent to using different training data for different heads in each iteration. The proposed DiverseHead differs from CCT, which does not distinguish the main or auxiliary decoder. Also, all decoders are trained by both supervised and unsupervised losses. In addition, the proposed method combines self-training and consistency regularization, which calculates unsupervised loss by using optimized pseudo labels. Specially, given are both a labelled dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing M images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with N images. The network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{len(heads)}$. Each of these heads is initialized differently. The expectation is to gain the trained network $Q$ leveraging these labelled and unlabelled data based on the proposed semi-supervised learning methods.
The output of each head refers to a probability-based prediction for all classifications corresponding to its input. When working with labelled data, the supervised loss $\mathcal{L}_{sup}^{j}$ between the $k\leavevmode\nobreak\ th$ ground truth $y_{k}$ and its corresponding prediction $p_{k}^{j}$ for $j\leavevmode\nobreak\ th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
$\mathcal{L}_{sup}^{j}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{j},y_{k}\right)\right),$ (1)
where $W$ and $H$ refer to the weight and height of input images. The ultimate supervised loss is determined by the mean of the losses across all heads
$\mathcal{L}_{sup}=\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}.$ (2)
In each iteration, we also get prediction $r_{k}^{j}$ corresponding to unlabelled data $u_{k}^{j}$ from $j\leavevmode\nobreak\ th$ head. In order to get higher precision pseudo labels, inspired by ensemble learning, the predictions from all heads are aggregated to create a combined prediction $\hat{r}_{k}^{sum}$. Subsequently, this combined prediction is utilized to calculate pseudo labels $\hat{y}_{k}^{sum}$ through an $argmax$ operation, which returns the indices of the maximum values of the prediction along the class dimension. Next, there is a random selector to pick only one head as the main branch to be supervised by minimizing the unsupervised loss
$\mathcal{L}_{unsup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(r_{k}^{main},\hat{y}_{k}^{sum}\right)\right),$ (3)
where $r_{k}^{main}$ is the prediction from the selected head.
Finally, the whole loss is written as
$\mathcal{L}=\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup},$ (4)
where $\lambda$ is the trade-off weight between the supervised and unsupervised losses.
Two perturbation methods. Based on the framework of DiverseHead, we propose a parameter perturbation method called dynamic freezing and also apply dropout as another perturbation to increase the diversity of features during training. Specifically, the first step of dynamic freezing refers to a process where, during each iteration, half of the heads are randomly selected to be frozen. This implies that the parameters within these selected heads will remain unchanged and not undergo updates. Every head has an equal probability of being chosen. Before going to the next iteration, those frozen heads are unfrozen. The algorithm of the DiverseHead semi-supervised learning approach with the perturbation of dynamic freezing is expressed in the Algorithm 1.
Another form of perturbation involves the utilization of dropout layers. We introduce a dropout layer before the last convolutional layer for each head. The pseudocode of this method is expressed in Algorithm 2. In the training phase, specific components of the feature tensor are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution. This dropout operation is employed to enhance the variability of the output predictions. In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.3. Similar to the training procedure of the semi-supervised learning approach based on dynamic freezing, the dropout-based approach adheres to the same training pipeline outlined above, with the distinction that all heads remain unfrozen. To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework. The performance of each combination is assessed in Section V.
Algorithm 1 DiverseHead Semi-Supervised Learning with Dynamic Freezing Pseudocode
  Initialization:
              Initialise backbone using ResNet-50
              Randomly initialise parameters for each head
  INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
               Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
[…]
        $\left\{head^{i}\right\}_{i=random}^{len(heads)/2}\leftarrow sample\left(\left\{head^{i}\right\}_{i=1}^{len(heads)}\right)$
        $Freeze(\left\{head^{i}\right\}_{i=random}^{len(heads)/2})$
        for $j\in\left\{1,...,len(heads)\right\}$ do
            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
            $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
        $r_{k}^{sum}=sum\left(\left\{r_{k}\right\}_{k=1}^{len(heads)}\right)$
        $\hat{y}_{k}^{sum}\leftarrow argmax(r_{k}^{sum})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{len(heads)}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{sum})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
        $Unfreeze(\left\{head^{i}\right\}_{i=random}^{len(heads)/2})$
  OUTPUT: Trained Model Q
Algorithm 2 DiverseHead Semi-Supervised Learning with Dropout Pseudocode
  Initialization:
          Initialise backbone using ResNet-50
          Randomly initialise parameters for each head
          Add dropout layer before the last convolutional layer
  INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
               Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
[…]
        for $j\in\left\{1,...,len(heads)\right\}$ do
            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
            $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
        $r_{k}^{sum}=sum\left(\left\{r_{k}\right\}_{k=1}^{len(heads)}\right)$
        $\hat{y}_{k}^{sum}\leftarrow argmax(r_{k}^{sum})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{len(heads)}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{sum})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
  OUTPUT: Trained Model Q
III Cross-model Supervision (DiverseModel)
Figure 3: DiverseModel: an online semi-supervised learning approach.
The proposed DiverseModel differs from CPS, exploring different networks in parallel to generate comprehensive pseudo labels. Since different networks pay different and complementary attention to the same input, this offers the basis that they are able to benefit from each other. DiverseModel structure includes three distinct networks which are PSPNet [zhao2017pyramid], UNet [ronneberger2015u], SegNet [badrinarayanan2017segnet], respectively. Algorithm 3 presents the pseudocode of the DiverseModel method.
The labelled data is used in a regular supervised learning manner to train these models by using the standard cross-entropy loss function $\ell_{ce}$. The supervised loss is expressed as:
$\mathcal{L}_{sup}=\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{n},y_{k}\right)\right),$ (5)
where $p_{k}^{n}$ represents the prediction of $k\leavevmode\nobreak\ th$ pixel from $n\leavevmode\nobreak\ th$ network.
In addition, unlabelled data is used to generate pseudo labels, which are then exploited for cross-supervision to inform each network. Different from the version presented in [ma2023confidence], all loss calculations in this work solely focus on cross-entropy loss to avoid the effects resulting from different types of loss functions on performance. The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through $argmax$ operation. For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from first network and the pseudo label $r^{2}$ generated by the second network is defined as:
$\mathcal{L}_{unsup}^{12}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{1},\hat{r}_{k}^{2}\right)\right).$ (6)
Algorithm 3 DiverseModel Semi-supervised Learning Pseudocode
  Initialization:
               Randomly initialise 3 models, PSPNet $Q^{1}$, UNet $Q^{2}$, SegNet $Q^{3}$
  INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
               Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
[…]
        $\mathcal{L}_{sup}=loss(Q^{1}(x_{k}),y_{k})$ + $loss(Q^{2}(x_{k}),y_{k})$
                  + $loss(Q^{3}(x_{k}),y_{k})$ based on (1)
        $\hat{r}_{k}^{1}\leftarrow argmax(Q^{3}(u_{k}))$
        $\hat{r}_{k}^{2}\leftarrow argmax(Q^{3}(u_{k}))$
        $\hat{r}_{k}^{3}\leftarrow argmax(Q^{3}(u_{k}))$
        $\mathcal{L}_{unsup}=loss(Q^{1}(u_{k}),\hat{r}_{k}^{2})$ + $loss(Q^{1}(u_{k}),\hat{r}_{k}^{3})$
                  + $loss(Q^{2}(u_{k}),\hat{r}_{k}^{1})$ + $loss(Q^{2}(u_{k}),\hat{r}_{k}^{3})$
                  + $loss(Q^{3}(u_{k}),\hat{r}_{k}^{1})$ + $loss(Q^{3}(u_{k}),\hat{r}_{k}^{2})$
        $\mathcal{L}\leftarrow\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
  OUTPUT: Trained Model $Q^{1}$, $Q^{2}$, $Q^{3}$
The cross-pseudo supervision among three networks creates six losses by using the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the result of linearly adding the six individual losses, as shown below
$\begin{array}[]{r}\mathcal{L}_{unsup}=\mathcal{L}_{unsup}^{12}+\mathcal{L}_{unsup}^{13}+\mathcal{L}_{unsup}^{21}\\ +\mathcal{L}_{unsup}^{23}+\mathcal{L}_{unsup}^{31}+\mathcal{L}_{unsup}^{32}.\end{array}$ (7)
The total loss $\mathcal{L}$ is the linear addition of $\mathcal{L}_{sup}$ and $\mathcal{L}_{unsup}$, which is previously given in equation (4)
IV Dataset Description
We employed diverse datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset [rottensteiner2012isprs], (2) the DFC2020 dataset [robinson2021global], (3) the RoadNet dataset [liu2018roadnet], and (4) the Massachusetts Buildings dataset [MnihThesis]. In the sequel, we share the details of each datasets utilised in this paper.
1) ISPRS Potsdam Semantic Labeling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS). The ground sampling distance of the two modalities of the true orthophoto (TOP) and the DSM is 5 cm. This dataset was annotated manually into six land cover classes, which are impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background. This dataset provides 38 multi-source patches (all with the size of $6000\times 6000$ pixels), which contain infrared (IR), red, green and blue bands orthorectified optical images with corresponding digital surface models (DSM). For computation purposes, we partitioned all these data tiles into $512\times 512$ patches, resulting in 3456, 201, and 1815 samples for training, validation, and test sets, respectively. We randomly select 1/4 of training data as labelled data and use the remaining 3/4 as unlabelled data for semi-supervised learning.
2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset which is based on the SEN12MS dataset [schmitt2019sen12ms]. It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps across seven areas namely Mexico City, Kippa-Ring, Khabarovsk, Black Forest, Mumbai, Cape Town, Bandar Anzali between 2016 and 2017. The size of all patches is $256\times 256$ pixels. There are 6112, 986 and 5127 images for training, validation and test sets, respectively. In this paper, we utilize only one-fifth of the labelled data for training, while the remaining four-fifths are employed as unlabeled data. The fine-grained International Geosphere-Biosphere Program (IGBP) classification scheme in SEN12MS was aggregated to ten coarser-grained classes, which are forest (11.3%), shrubland (6.9%), savanna (23.6%), grassland (16.8%), wetlands (1.1%), croplands (17.9%), urban/built-up (10.6%), snow/ice (0.0%), barren (5.2%), and water (6.5%).
3) RoadNet is a benchmark dataset for road detection with 0.21-m spatial resolution per pixel covering 21 typical urban regions about 8 $km^{2}$ in Ottawa, Canada. The annotation data includes road surfaces, road edges, and road centerlines for each image. we only used road surfaces for the segmentation task. the number of samples for training, validation and testing is 410, 45 and 387, respectively. In the semi-supervised learning approaches, only 1/4 of the annotated training data is used, while the remaining 3/4 of the training data are employed as unlabeled data for training purposes.
4) Massachusetts Buildings Dataset predominantly encompasses urban and suburban regions, encompassing structures of varying scales. This includes diverse buildings such as individual houses and garages, all of which are included in the labels. The dataset consists of 151 aerial images capturing the Boston area covering roughly 340 square kilometres, with a resolution of $1\leavevmode\nobreak\ m^{2}$ per pixel. The dataset was divided into subsets as follows: 137 images for training, 10 images for testing, and 4 images for validation. Each subset consists of images sized $1500\times 1500$ pixels. One-quarter of the annotated training images and the remaining three-quarters of the unlabelled training data are utilized in semi-supervised learning.
V experiment and results analysis
V-A Implementation Details
We implemented our methods by using the PyTorch framework. Following [chen2018encoder], we used a mini-batch SGD optimizer and adopted a poly learning rate policy, where the current learning rate equals the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$. The initial learning rate and power are set to 0.01 and 0.9, respectively. All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  [gw4].
We implemented various classic semi-supervised learning methods (MT, CCT, GCT and CPS) on remote sensing datasets. All these methods use DeepLabv3+ with ResNet-50 pre-trained on ImageNet as the semantic segmentation network. To be fair the proposed DiverseHead is also based on DeepLabv3+ with the pretrained ResNet-50.
We comprehensively analyzed all methods by quantifying performance via class-related measures, including overall accuracy (OA), user’s accuracy (UA), producer’s accuracy (PA), mean intersection over union (mIoU), and $F_{1}$-score. Expressions of all five performance metrics are given as follows: $\mathrm{OA}=\frac{TP+TN}{TP+TN+FP+FN}$, $\mathrm{UA}=\frac{TP}{TP+FP}$, $\mathrm{PA}=\frac{TP}{TP+FN}$, $\mathrm{mIoU}=\frac{|\mathrm{TP}|}{|\mathrm{TP}+\mathrm{FN}+\mathrm{FP}|}$, $F_{1}=\frac{2\cdot PA\cdot UA}{PA+UA}$, where TP, TN, FP, and FN refer to the numbers of pixels that are true positives, true negatives, false positives, and false negatives for each class, respectively.
V-B Quantitative Results and Analysis
We evaluated the performance of the proposed approach and reference networks using 4 aforementioned datasets, and the average results are presented in Table I. The reason we first share average results across datasets is to provide a global applicability display for all utilised techniques, and each data set’s performance will be discussed in the sequel. Evaluating the performance results in Table I, we conclude that the proposed DiverseModel achieves the best performance among all metrics. In particular, for metrics like UA and mIoU, the proposed DiverseModel exhibits an improvement of over 5% compared to another cross-model-based approach, CPS. Although the performance of DiverseHead is lower than that of DiverseModel, it outperforms other reference methods and reaches the second best. Upon convergence of contemporary methodologies with remote sensing datasets, MT demonstrates notably inferior average performance across various metrics, with the exception of UA. Both CCT and GCT exhibit comparable performance, yet the performance superiority of CCT is notably evident in its PA metric. CPS consistently yields superior outcomes across the majority of metrics when juxtaposed against other aforementioned state-of-the-art methodologies.
Meanwhile, it is worth noting that the DiverseHead consists of a notably lighter architecture compared to all other models where Table II presents the required number of parameters during training for each semi-supervised learning architecture. It is evident that the architectures using multiple segmentation networks during training typically need higher memory demands. In contrast, the proposed DiverseHead is remarkably lightweight as it only employs a single model with multiple heads which consists of only 2 convolution layers, eliminating the need for multiple models during training. Thus, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table II), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network. Despite this need for a lower number of parameters, the DiverseHead proposed method outperforms all state-of-the-art with at least 1% accuracy whilst obtaining similar performance to the companion method of DiverseModel.
Table I: Average performance comparison with the state of the arts on 4 datasets. DT and DF present Dropout and dynamic freezing, respectively.
OA UA PA mIoU $F_{1}$
MT[tarvainen2017mean] 85.63% 73.14% 80.10% 63.93% 76.39%
CCT[ouali2020semi] 86.60% 72.17% 83.29% 64.07% 77.22%
GCT[ke2020guided] 86.53% 73.33% 80.50% 64.59% 76.72%
CPS[chen2021semi] 87.17% 72.89% 84.62% 65.02% 78.18%
DiverseHead (DT) 88.42% 77.80% 84.95% 69.70% 81.05%
DiverseHead (DF) 88.29% 76.98% 84.87% 68.85% 80.58%
DiverseModel 88.86% 78.24% 86.06% 70.53% 81.79%
Table II: The required parameter size in each semi-supervised learning approach. Base means supervised segmentation network DeepLabv3+. DT and DF present Dropout and dynamic freezing, respectively.
CPS [chen2021semi] MT [tarvainen2017mean] CCT [ouali2020semi] GCT [ke2020guided]
Size (MB) 303.378 303.378 337.655 335.049
Base DiverseHead(DT) DiverseHead(DF) DiverseModel
Size (MB) 151.689 175.806 175.806 936.262
Specifically, we present the results of each semi-supervised learning method for the 4 utilised datasets in Table III. DiverseNet architectures reach a remarkable performance for each of the datasets. Specifically, the proposed DiverseModel achieves the best performance on both RoadNet and Massachusetts building datasets for all metrics, and the DiverseHead (DT) shows the second best for the most of metrics on these two datasets. For the other two datasets, Potsdam and DFC2020, The DiverseHead (DF) method shows competitive and better performance compared to the companion DiverseModel for most of the metrics.
Table III: Performance comparison with the state of the arts on 4 datasets. DT and DF present Dropout and dynamic freezing, respectively.
OA UA PA mIoU $F_{1}$
Potsdam MT[tarvainen2017mean] 81.98% 73.66% 78.39% 63.07% 75.95%
CCT[ouali2020semi] 82.66% 74.64% 77.62% 64.16% 76.10%
GCT[ke2020guided] 83.99% 75.65% 80.81% 65.80% 78.14%
CPS[chen2021semi] 85.00% 75.76% 82.94% 66.69% 79.19%
DiverseHead (DT) 84.60% 77.42% 80.49% 67.18% 78.92%
DiverseHead (DF) 85.37% 78.43% 81.86% 68.49% 80.11%
DiverseModel 85.76% 76.75% 83.45% 67.85% 79.96%
DFC2020 MT[tarvainen2017mean] 78.64% 59.59% 73.57% 50.44% 65.85%
CCT[ouali2020semi] 79.71% 59.87% 76.40% 51.04% 67.13%
GCT[ke2020guided] 80.84% 61.47% 71.43% 52.17% 66.07%
CPS[chen2021semi] 81.50% 61.75% 79.47% 53.21% 69.50%
DiverseHead (DT) 82.11% 62.31% 80.38% 53.93% 70.20%
DiverseHead (DF) 82.19% 62.29% 80.73% 54.02% 70.32%
DiverseModel 81.87% 62.20% 80.69% 53.71% 70.25%
RoadNet MT[tarvainen2017mean] 93.00% 80.74% 85.54% 72.96% 83.07%
CCT[ouali2020semi] 93.90% 83.85% 87.26% 76.23% 85.52%
GCT[ke2020guided] 93.57% 79.26% 89.25% 73.51% 83.96%
CPS[chen2021semi] 94.36% 84.21% 88.85% 77.44% 86.47%
DiverseHead (DT) 96.28% 90.89% 91.95% 84.84% 91.42%
DiverseHead (DF) 95.36% 87.99% 90.31% 81.35% 89.13%
DiverseModel 96.79% 92.82% 92.57% 86.87% 92.70%
Massachusetts MT[tarvainen2017mean] 88.92% 78.56% 82.90% 69.25% 80.67%
CCT[ouali2020semi] 89.65% 79.91% 84.23% 70.95% 82.01%
GCT[ke2020guided] 87.73% 76.95% 80.51% 66.88% 78.69%
CPS[chen2021semi] 87.85% 69.85% 87.24% 62.76% 77.58%
DiverseHead (DT) 90.69% 80.56% 86.97% 72.86% 83.64%
DiverseHead (DF) 90.25% 79.22% 86.59% 71.53% 82.74%
DiverseModel 91.01% 81.19% 87.53% 73.67% 84.24%
We also considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS). The segmentation network consists of only a single head. During each iteration, the current model’s predictions for unlabeled data are utilized to generate pseudo labels through an $\arg\max$ operation. These generated pseudo labels are then used to calculate the unsupervised loss. The baseline is the segmentation network only trained by labelled data. The performance of the baseline and SHS, along with the proposed DiverseHead, are presented in Table IV in order to provide experimental evidence of the novelty of utilisation of multiple heads in DiverseHead architecture. While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the baseline model especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads. Specifically, the mIoU of DiverseHead (DF) is 5.7% higher than that of Baseline and approximately 4% higher than that of SHS for Potsdam. Also, the mIoU value of DiverseHead (DT) is about 10% higher than that of SHS.
Table IV: performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present Dropout and dynamic freezing, respectively.
OA UA PA mIoU $F_{1}$
Potsdam Baseline 81.64% 73.86% 76.05% 62.72% 74.94%
SHS 83.43% 74.57% 79.93% 64.51% 77.16%
DiverseHead (DT) 84.60% 77.42% 80.49% 67.18% 78.92%
DiverseHead (DF) 85.37% 78.43% 81.86% 68.49% 80.11%
RoadNet Baseline 93.56% 82.85% 86.53% 75.06% 84.65%
SHS 93.79% 81.19% 88.62% 74.87% 84.74%
DiverseHead (DT) 96.28% 90.89% 91.95% 84.84% 91.42%
DiverseHead (DF) 95.36% 87.99% 90.31% 81.35% 89.13%
Table V: Ablation study about the number of heads for DiverseHead with dynamic freezing
# of heads OA UA PA mIoU $F_{1}$
Potsdam 8 85.70% 78.47% 82.09% 68.82% 80.24%
10 85.37% 78.43% 81.86% 68.49% 80.11%
12 85.51% 77.23% 83.03% 68.14% 80.02%
RoadNet 8 94.28% 80.20% 92.23% 75.59% 85.79%
10 95.36% 87.99% 90.31% 81.35% 89.13%
12 95.35% 85.57% 92.22% 80.52% 88.77%
Average 8 89.99% 79.33% 87.16% 72.21% 83.02%
10 90.36% 83.21% 86.08% 74.92% 84.62%
12 90.43% 81.40% 87.62% 74.33% 84.40%
An ablation study was conducted on the Potsdam and RoadNet datasets to explore the impact of the number of heads for the proposed DiverseHead method with dynamic freezing perturbation. The results in Table V indicate that using 10 heads yields the optimal choice across most metrics in terms of average performance. The performance difference between the versions with different numbers of heads is very small on Potsdam, but the version of 8 heads shows a performance drop. Considering the stability, We chose the 10-head version of DiverseHead (DF) for all experiments in this paper.
Table VI: Performance comparison of the DiverseModel with its constituent networks
OA UA PA mIoU $F_{1}$
UNet 83.28% 74.18% 78.88% 64.51% 76.46%
SegNet 82.37% 73.53% 77.46% 63.24% 75.45%
PSPNet 81.96% 74.23% 76.99% 63.08% 75.58%
DiverseModel 85.76% 76.75% 83.45% 67.85% 79.96%
Since the DiverseModel technique uses three different segmentation networks, we also evaluated the performance of each individual member network within the DiverseModel in Table VI. The results indicate that the cross-supervision of all networks shows a significant improvement in the segmentation task performance compared to each separate network.
Figure 4: Visual results of each method on Potsdam Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
Figure 5: Visual results of each method on RoadNet Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
Figure 6: Visual results of each method on DFC2020 Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
Figure 7: Visual results of each method on Massachusetts Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
V-C Qualitative Results and Analysis
The visible results are presented in Figures from 4 to 7. For each dataset, we chose one case and depicted the RGB image, ground truth, and predictions of all the methods. Compared to the ground truth, the proposed three methods (right-top 3 sub-figures) achieve better similarity to the ground truth compared to the reference models. We also calculated the IoU scores for all predictions with the ground truth for this case, which are shown in the sub-caption of each prediction. The IoU scores also demonstrate that the performance of the proposed DiverseNet methods is more competitive than that of the other approaches.
VI Conclusion
In this paper, we proposed a lightweight semi-supervised learning based on a multi-head structure called DiverseHead, which equipped a dynamic freezing and dropout perturbation approaches. The multi-head idea benefits from bagging and shows competitive performance for the segmentation of remote sensing imagery. Also, another high-efficiency multiple network based semi-supervised learning method called DiverseModel is investigated for further analysis. DiverModel shows the strongest performance on various datasets compared to the state of the arts but requires more computational resources, whilst the performance of DiverseHead is slightly lower than DiverseModel but dramatically lighter than all reference models.