### v2

# DiverseNet: Decision Diversified Semi-supervised Semantic Segmentation Networks for Remote Sensing Imagery

Abstract
Semi-supervised learning aims to help reduce the cost of the manual labelling process by leveraging valuable features extracted from a substantial pool of unlabeled data alongside a limited set of labelled data during the training phase. Since pixel-level manual labelling in large-scale remote sensing imagery is expensive, semi-supervised learning becomes an appropriate solution to this. However, most of the existing consistency learning frameworks based on network perturbation are very bulky. There is still a lack of lightweight and efficient perturbation methods to promote the diversity of features and the precision of pseudo labels during training. In order to fill this gap, we propose DiverseNet which explores multi-head and multi-model semi-supervised learning algorithms by simultaneously enhancing precision and diversity during training. The two proposed methods in the DiverseNet family, namely DiverseHead and DiverseModel, both achieve the better semantic segmentation performance in four widely utilised remote sensing imagery data sets compared to state-of-the-art semi-supervised learning methods. Meanwhile, the proposed DiverseHead architecture is simple and relatively lightweight in terms of parameter space compared to the state-of-the-art methods whilst reaching high-performance results for all the tested data sets.
keywords:
Semi-supervised Learning, Consistency Learning, Semantic Segmentation
[…]
organization=School of Computer Science and Informatics, Cardiff University,city=Cardiff, postcode=CF24 4HJ, country=United Kingdom
1 Introduction
Supervised deep learning has gradually become the dominant technique in computer vision, especially during the last decade. Specifically, with its great success in semantic segmentation in computer vision, automated land cover classification approaches have been significantly improved  (Vali et al., 2020). Nevertheless, supervised learning necessitates a substantial and meticulously labelled dataset. In the case of extensive remote sensing data, such as satellite imagery and drone-captured images in complex terrains, acquiring pixel-wise expert annotations is a time-consuming, labour-intensive, and costly endeavour. While the field of computer vision provides numerous well-annotated datasets, transferring deep learning models trained on these datasets to the remote sensing domain is a formidable challenge. This is primarily due to the significant differences between images generally used in computer vision and remote sensing data, including hyperspectral and synthetic aperture radar (SAR) imagery, in which images often possess unconventional and non-intuitive characteristics. Fortunately, thanks to open-access remote sensing data sources, a large amount of unlabelled imagery can be accessed easily and freely. Thus, only using a limited amount of labelled data and exploiting the benefits of a large amount of unlabelled data could be a feasible solution for solving the problem of the lack of labelled remote sensing datasets (Wang et al., 2022a; Yang et al., 2023; Wang et al., 2023).
In computer vision, semi-supervised learning has become a popular research direction to eliminate the labour-intensive and expensive annotation stages mentioned above. It leads to competitive performance compared to supervised learning in applications such as image classification (Zhang et al., 2021; Huang et al., 2021; Wang et al., 2021b) and segmentation (Hu et al., 2021; Wang et al., 2022b, 2021a), which exploit just a part of labelled data and a large amount of unlabelled data. Specifically, mainstream semi-supervised learning methods exploit “pseudo” labels that are fake and/or auto-generated label information obtained as a result of an unsupervised step. Since pseudo labels are involved in the model training stage, the accuracy of pseudo labels is regarded as a key factor affecting the performance of semi-supervised learning approaches. However, the accuracy of pseudo labels is not always satisfactory especially when labelled training data is not complete and sufficient enough. Meanwhile, increasing diversity of the pseudo-labels has become another key research direction within the semi-supervised learning domain, especially for consistency regularization (French et al., 2019). This is because a greater diversity of pseudo-labels enhances the effectiveness of training deep learning networks.
In order to empower networks to harness the potentials within unlabeled data, a notable technique known as consistency regularization has emerged as a widely adopted method for semi-supervised learning. (French et al., 2019; Filipiak et al., 2021). Specifically, consistency regularization methods are built up on the theory of assumption of smoothness, which posits that if two points reside in a high-density region of feature space and are close to each other, their corresponding labels should be the same or consistent (Chen and Wang, 2010; Luo et al., 2018; French et al., 2019). In practice, these methods for consistency regularization ensure adherence to this assumption by compelling networks to generate predictions that remain consistent across modified iterations of unlabeled inputs.
The strategies for the aforementioned perturbations can be categorised into three groups, namely input, feature, and network perturbations. Input perturbation is a method that involves making adjustments or alterations to the input data as part of the training process. In this way, the semi-supervised learning approach promotes the consistency of predictions made in response to these modified inputs. Widely used input perturbation methods employed in semi-supervised learning include Cutout (DeVries and Taylor, 2017), MixUp (Zhang et al., 2017), and CutMix (Yun et al., 2019). Among them, CutMix stands out by combining aspects of MixUp and CutOut techniques, demonstrating superior performance compared to them (French et al., 2019). However, input perturbation introduces artificial noise to the data, which might lead to incorrect or noisy labels when generating pseudo-labels for unlabeled examples. This might negatively impact the training process by providing incorrect guidance to the model when using pseudo labels. Cross-consistency training (CCT) (Ouali et al., 2020) presents a feature perturbation method that introduces noise to both low- and high-level features. These perturbed features are then fed into multiple decoders to generate multiple outputs, followed by the enforcement of consistency among the outputs obtained from different decoders. However, feature perturbation alters the intrinsic characteristics and relationships between features, which also might lead to the creation of artificial feature representations that do not accurately reflect the underlying patterns in the original data. Consequently, the perturbed features could result in incorrect or noisy labels when generating pseudo-labels for unlabeled examples, which negatively impacts the learning process. Lastly, network perturbation uses different networks or the same network with different initialization weights to promote diversity of predictions, such as cross pseudo supervision (CPS) (Chen et al., 2021), confidence-guided semi-supervised learning (CGSSL) (Ma et al., 2023), mean teacher (MT), (Tarvainen and Valpola, 2017), CCT (Ouali et al., 2020), guided collaborative training (GCT) (Ke et al., 2020) and cross teaching between CNN and transformer (Luo et al., 2022). Compared to the previous two perturbation methods, although network perturbation methods introduce perturbations created by the model itself instead of artificial noise, they generally require rapidly increased computational resources due to the increased number of networks (whole or their inner stages) that are used.
(a) DiverseHead
(b) DiverseModel
[…]
1.1 Contributions
The preceding overview underscores that, although previous perturbation techniques in semi-supervised learning have their advantages, they also bring inherent drawbacks which could be a key consideration affecting the performance in semi-supervised learning. In addressing the challenges posed by perturbation-based semi-supervised learning and its complexity, it becomes imperative to explore more sophisticated and efficient approaches. These approaches aim to enhance the capabilities of semi-supervised learning frameworks by integrating advanced perturbation methods. This paper proposes two perturbation-based semi-supervised network architectures, coined as DiverseNet, which consist of decision (DiverseHead) and novel feature (DiverseModel) diversifying network architectures for the purposes of semi-supervised semantic segmentation of remote sensing imagery. The proposed lightweight semi-supervised semantic segmentation network called DiverseHead is shown in Figure 1-(a) based on multiple decision heads to promote precision and diversity of pseudo labels whilst using two novel perturbation methods of dropout (DT) and dynamic freezing (DF). In addition, we also analysed a previously proposed high-performance multi-model structure called DiverseModel (Ma et al., 2023) as shown in Figure 1-(b) for scenarios equipped with high-memory computational resources. Specifically, the contributions of this work are as follows:
  1. 1.
    We proposed a simple but efficient semi-supervised learning architecture, promoting the utilisation of multiple decision heads called DiverseHead. This structure is inspired by bagging (also called bootstrap aggregating) and provides diversity for parameters and features. Since each head only consists of a couple of convolutional layers instead of a whole network, the architecture is relatively lightweight and adaptable to various segmentation networks.
  2. 2.
    In order to provide efficient pseudo labels in the training stage, we provide a comprehensive voting mechanism based on the proposed DiverseHead structure. The voting mechanism not only uses the joint output from multiple heads but also individually considers pseudo labels generated from these outputs, which are called Mean Voting and Max Voting in this paper, respectively.
  3. 3.
    For the purposes of creating perturbation in diversifying decisions, we used two important techniques, namely ‘dynamic freezing’ and ‘dropout’. We make these techniques fit with the DiverseHead semi-supervised learning architecture to diversify the parameters of the network. The two perturbation methods differ from input/feature perturbation methods since the source of feature diversity is from parameter diversity instead of introducing artificial noise to data or features.
  4. 4.
    Further analysis and more detailed evaluation has been carried out on a previously proposed architecture DiverseModel (Ma et al., 2023) which is based on consistency learning among multiple neural networks with distinct weights. We provide a detailed comparison study for various semantic segmentation datasets in this paper. Also, we use Grad-CAM (Selvaraju et al., 2020) to verify the observation that different networks exhibit varied attention to the same input, culminating in the combined pseudo label derived from the multiple model predictions showcasing superior quality.
The rest of the paper is organised as follows: Section 2 discusses the related work on semi-supervised semantic segmentation in remote sensing and some basic knowledge on ensemble machine learning whilst in Section 3 and 4, the proposed algorithms DiverseHead and DiverseModel are presented. Section 5 describes the utilised segmentation dataset of remote sensing imagery. The experimental setting and the qualitative and quantitative analyses of the experimental results are performed in Section 6. Section 7 concludes the paper with a summary.
2 Related Work
Semantic segmentation is a well-studied and wide-spread research topic in remote sensing, incorporated within applications such as land cover classification/mapping (Dong et al., 2019), building change detection (Zheng et al., 2022), road extraction (Ghandorh et al., 2022), and marine debris detection (Kikaki et al., 2022; Booth et al., 2023). The volume of research on semantic segmentation in remote sensing in the literature is rapidly increasing with the success of deep learning in computer vision, due to the strong task-similarity between two areas. Fully convolutional networks (FCNs) (Long et al., 2015) is one group of the most widespread networks for segmentation tasks. They have made a considerable contribution to various segmentation tasks either in remote sensing or computer vision. Following the FCNs’ success, SegNet (Badrinarayanan et al., 2017) and UNet (Ronneberger et al., 2015) adopt a symmetrical encoder-decoder structure with skip connections, leveraging multi-stage features within the encoder. Alternatively, PSPNet (Zhao et al., 2017) introduces a pyramid pooling structure that helps provide a global contextual understanding for pixel-level scene parsing. DeepLab architecture (Chen et al., 2017) proposes a new type of convolution called atrous convolution and an atrous spatial pyramid pooling (ASPP) operation to enable the network to have the ability to control the spatial perception range of convolution kernels by setting different dilation rates. DeepLab has been extended to DeepLabv3+ (Chen et al., 2018) with an improved hybrid approach that combines an encoder-decoder structure and the ASPP. Within DeepLabv3+, the ASPP module is incorporated following a ResNet backbone, enabling the network to delve into deeper layers and extract high-level features, which is helpful for accurately segmenting pixels around boundaries (Xia et al., 2021; Zhang et al., 2019). DeepLabv3+ is one of the most used networks in the literature for semi-supervised learning segmentation in the computer vision area.
Semi-supervised learning aims to alleviate the need for expensive annotation work and improve the models’ performances by making use of unlabelled data. Self-training (Lee et al., 2013) (also known as pseudo labelling) represents one of the primitive semi-supervised learning strategies for both classifications (Kim et al., 2020) and segmentation (Zhu et al., 2021), and involves the generation of pseudo labels through model predictions. These pseudo-labels are subsequently used to retrain the model. Various variations of self-training have been proposed focusing on determining the appropriate pseudo labels. Another widely developed semi-supervised learning approach called consistency regularization (French et al., 2019) is to force networks to give consistent predictions for unlabeled inputs that undergo diverse perturbations. CCT (Ouali et al., 2020) employs an encoder-decoder architecture with multiple auxiliary decoders. These decoders introduce diversity in the output by feature perturbations specific to each auxiliary decoder. They calculate MSE loss between predictions of the main decoder and each auxiliary decoder without creating pseudo labels. It is worth noting that the unsupervised loss is not used to supervise the main decoder. Following CCT, subsequent methods like GCT (Ke et al., 2020) and CPS (Chen et al., 2021) have been proposed to introduce network perturbation for consistency regularization. Specifically, they use two same network structures but with different weight initialization. CPS differs from GCT by employing pseudo labels to enforce consistency, whereas GCT achieves consistency regularization by utilizing predictions of networks cooperating with a flow detector. Although the network perturbation was achieved through different weight initialization in GCT and CPS, the ability to generate diversity is still limited. ICNet (Wang et al., 2022a) was proposed to use teacher networks for increasing the model difference to enhance the diversity of pseudo labels based on an iterative contrast network. Nonetheless, these methods consistently rely on the same network architecture. To introduce more diversity, we advocate the use of different model architectures. By leveraging various networks, we can obtain distinct and complementary features from these models even with the same input data. Thus, the proposed DiverseModel differs from the aforementioned methods, exploring different networks in parallel to generate more diversity of pseudo labels to improve training effectiveness.
Ensemble machine learning is a concept that employs multiple learners and combines their predictions (Sewell, 2008). Bagging, a form of ensemble learning, is a technique aimed at reducing prediction variance by creating multiple iterations of a predictor and then utilizing them to form an aggregated predictor (Breiman, 1996). Specifically, bagging creates sample subsets by randomly selecting from the training data set and subsequently utilises these acquired subsets to train the foundational models for integration. When forecasting a numerical result, the aggregation involves taking an average of the versions, while for predicting a class it relies on a majority vote. Bagging is a commonly employed approach for enhancing the robustness and precision of machine learning algorithms for classification and regression (Ren et al., 2016).
3 DiverseHead (Cross-head Supervision)
Figure 2: DiverseHead: an online semi-supervised learning approach. This figure applies the dynamic freezing strategy: slide windows of the freezer randomly select a certain number of heads to freeze the parameter of heads (not updated by backpropagation). Additionally, during every iteration, all heads undergo supervision through a supervised loss, yet only a single head is randomly chosen to be updated by unsupervised loss.
Figure 3: The Proposed Voting Module: a voting mechanism for the pseudo label creation. In the unsupervised part, the voting module combines the mean output of multiple heads (mean voting) and individual pseudo labels (max voting) to generate more efficient pseudo labels. Argmax returns the indices of the maximum values of the prediction along the class dimension. The dashed arrow serves as an illustration of a pixel voting for its classification in a segmentation map.
This section introduces the details of the proposed DiverseHead semi-supervised learning method which is proposed to create various predictions by multiple heads illustrated in Figure 2, specifically the number of convolutional layers for each head is 2. The source of perturbation is from two strategies, called dynamic freezing and dropout which are shown in Figure 4 and explained in detail in Section 3.1. The proposed DiverseHead does not distinguish between the main or auxiliary decoder, which differs from CCT. During every training iteration, all heads undergo supervision through supervised losses, yet only a single head is randomly chosen to be updated by an unsupervised loss. In addition, the proposed method combines self-training and consistency regularization, which uses their predictions to supervise itself and forces all perturbed head to give a consistent output. It makes an efficient voting module shown in Figure 3 to help calculate unsupervised loss by using optimized pseudo labels. To specifically explain the proposed semi-supervised framework, given both a labelled dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing $M$ images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with $N$ images, the network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{L}$, where L is the number of heads. Each of these heads is initialized differently. The expectation of the proposed semi-supervised learning architecture is to gain the trained network $Q$ leveraging these labelled and unlabelled data.
The output of each head refers to a probability-based prediction for all classifications corresponding to its input. When working with labelled data, the supervised loss $\mathcal{L}_{sup}^{j}$ between the $k^{\prime}th$ ground truth $y_{k}$ and its corresponding prediction $p_{k}^{j}$ for $j^{\prime}th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
$\mathcal{L}_{sup}^{j}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{j},y_{k}\right),$ (1)
where $W$ and $H$ refer to the weight and height of input images. The final supervised loss is determined by the mean of the losses across all heads
$\mathcal{L}_{sup}=\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}$ (2)
where L is the number of heads.
In each iteration, we also get prediction $r_{k}^{j}$ corresponding to unlabelled data $u_{k}^{j}$ from the $j^{\prime}th$ head. Inspired by ensemble learning, a voting module is proposed to get high-precision pseudo labels. There are two voting mechanisms in the proposed voting module, called mean voting and max voting, respectively. The former aggregates the predictions from all heads to create a combined prediction $\hat{r}_{k}^{mean}$. Subsequently, this combined prediction is used to calculate the mean pseudo labels $\hat{y}_{k}^{mean}$ through an $argmax$ operation, which returns the indices of the maximum values of the prediction along the class dimension. Apart from the mean voting module, the individual pseudo label is generated from the output of each head by the $argmax$ function. The latter regards all pseudo labels as voters, in which the mean pseudo label contributes $\varphi$ weight and each individual pseudo label contributes unit weight. $\varphi$ is a learnable parameter and its value changes depending on the dataset and training processing. The max voting module returns the class number that gets the most votes for each pixel. After max voting, an optimal pseudo label $\hat{y}_{k}^{final}$ is created to calculate unsupervised loss by using the cross-entropy loss function:
$\mathcal{L}_{unsup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(r_{k}^{main},\hat{y}_{k}^{final}\right),$ (3)
where $r_{k}^{main}$ is the prediction from the selected head. There is a random selector to pick only one head as the main branch to be supervised by minimizing the unsupervised loss.
Finally, the whole loss is written as
$\mathcal{L}=\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup},$ (4)
where $\lambda$ is the trade-off weight between the supervised and unsupervised losses.
(a) Dynamic Freezing
(b) Dropout
Figure 4: The proposed two perturbation methods: dropout and dynamic freezing. In (a), the dynamic selector is designed to amplify the parameter diversity of multiple heads. It randomly chooses a set number of heads not to be updated by backpropagation during each training iteration, thereby enhancing the variability of parameters.
3.1 Perturbation Methods
Algorithm 1 DiverseHead Semi-Supervised Learning with Dynamic Freezing Pseudocode
  Initialization:
              Initialise backbone using ResNet-50
              Initialise model $Q$
              Randomly initialise parameters for each head
  INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
               Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
  $L=length(heads)$
  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
        $\mathcal{R}=Randint(0,L,\frac{1}{2}L)$
        $Freeze(\left\{head^{i}\right\}_{i\in\mathcal{R}})$
        for $j\in\left\{1,...,L\right\}$ do
            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
            $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
            $\hat{y}_{k}^{j}\leftarrow argmax(r_{k}^{j})$
        $r_{k}^{mean}=sum\left(\left\{r_{k}^{j}\right\}_{j=1}^{L}\right)$
        $\hat{y}_{k}^{mean}\leftarrow argmax(r_{k}^{mean})$
        $\hat{y}_{k}^{final}\leftarrow voting(\left\{\hat{y}_{k}^{j}\right\}_{j=1}^{L},\hat{y}_{k}^{mean})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{L}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
        $\mbox{\it Unfreeze}(\left\{head^{i}\right\}_{i=R})$
  OUTPUT: Trained model $Q$
Based on the framework of DiverseHead, we propose a parameter perturbation method called dynamic freezing as shown in Figure 4 (a). The pseudocode of the algorithm for DiverseHead with the perturbation of dynamic freezing is given in Algorithm 1. Specifically, we use DeepLabv3+ with a ResNet-50 backbone as the segmentation network in our framework, ensuring consistency with the methods compared in this paper. The backbone is pretrained on ImageNet, and other parameters are initialised by Kaiming Initialization. Both labelled and unlabelled data are input to the framework, then the first step of dynamic freezing refers to a process where half of the heads are randomly selected to be frozen during each iteration. This implies that the parameters within these selected heads will remain unchanged and not undergo updates for the current iteration. Every head has an equal probability of being chosen. Following the cross-head supervision above, the segmentation model is updated by the supervised and unsupervised losses. Before going to the next iteration, those frozen heads are unfrozen.
Algorithm 2 DiverseHead Semi-Supervised Learning with Dropout Pseudocode
  Initialization:
          Initialise backbone using ResNet-50
          Randomly initialise parameters for each head
          Add dropout layer before the last convolutional layer
  INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
               Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
  $L=length(heads)$
  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
        for $j\in\left\{1,...,L\right\}$ do
            $p_{k}^{j}=Q^{head_{j}}(x_{k}^{j})$
            $\mathcal{L}_{sup}^{j}=loss(p_{k}^{j},y_{k})$ based on (1)
            $r_{k}^{j}=Q^{head_{j}}(u_{k}^{j})$
            $\hat{y}_{k}^{j}\leftarrow argmax(r_{k}^{j})$
        $r_{k}^{sum}=sum\left(\left\{r_{k}^{j}\right\}_{j=1}^{L}\right)$
        $\hat{y}_{k}^{mean}\leftarrow argmax(r_{k}^{mean})$
        $\hat{y}_{k}^{final}\leftarrow voting(\left\{\hat{y}_{k}^{j}\right\}_{j=1}^{L},\hat{y}_{k}^{mean})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{L}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
  OUTPUT: Trained Model Q
Another form of perturbation involves the use of dropout layers to the proposed DiverseHead structure to increase the diversity of features during training. We introduce a dropout layer after each convolutional block in each segmentation head of the network as shown in Figure 4 (b). The pseudocode of this method is given in Algorithm 2. As both dynamic freezing and dropout perturbation utilize the same semi-supervised learning framework, DiverseHead, the network initialization and supervised loss calculation procedures stay the same. Using dropout in DiverseHead, specific components of the weights in the heads of the network are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution. This dropout operation is employed to enhance the variability of the output predictions. In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.3. Although the dropout-based approach adheres to the same training pipeline with dynamic freezing outlined above, the distinction is that all heads remain unfrozen.
To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework. The performance of each combination is assessed in Section 6. Considering various datasets may require differing levels of diversity in pseudo labels within the proposed DiverseHead framework, adjustments can be made by altering the number of frozen heads and dropout rates in the methods of dynamic freezing and dropout, respectively.
4 Cross-model Supervision (DiverseModel)
Figure 5: DiverseModel: an online semi-supervised learning approach.
The proposed DiverseModel differs from CPS, exploring different networks in parallel to generate comprehensive pseudo labels. As shown in Figure 5, the DiverseModel structure includes three distinct networks, which can be various semantic segmentation networks. In this paper, we choose three widely used segmentation networks for experiments, which are PSPNet (Zhao et al., 2017), UNet (Ronneberger et al., 2015), SegNet (Badrinarayanan et al., 2017). Algorithm 3 presents the pseudocode of the DiverseModel method. Since different networks pay different and complementary attention to the same input, this offers the basis that they are able to benefit from each other. In order to provide evidence for this claim, we executed the Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2020) technique for every network employed within the framework of the DiverseModel architecture. Grad-CAM visualizes the areas of an image that are important to the model predictions from each network. Figure 6 depicts an example grad-CAM analysis for the BUILDING class in the Potsdam data set. Examining Figure 6, we can see that different networks pay different and complementary attention to the same input, and the pseudo labels from the DiverseModel prediction show the highest quality.
Figure 6: The upper section displays Grad-CAM outputs from individual networks within the DiverseModel architecture using the Potsdam dataset. The lower section showcases both the ground truth and pseudo labels generated through predictions from each network. In the lower right corner, the pseudo label from DiverseModel is presented.
The labelled data is used in a regular supervised learning manner to train these models by using the standard cross-entropy loss function $\ell_{ce}$. The supervised loss is expressed as:
$\mathcal{L}_{sup}=\frac{1}{3}\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{n},y_{k}\right),$ (5)
where $p_{k}^{n}$ represents the prediction of the $k\leavevmode\nobreak\ th$ pixel from the $n\leavevmode\nobreak\ th$ network.
In addition, unlabelled data is used to generate pseudo labels, which are then exploited for cross-supervision to inform each network. Different from the version presented in (Ma et al., 2023), all loss calculations in this work solely focus on cross-entropy loss to avoid the variations in performance resulting from different types of loss functions. The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through the $argmax$ operation. For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from the first network and the pseudo label $r^{2}$ generated by the second network is defined as:
$\mathcal{L}_{unsup}^{12}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{1},\hat{r}_{k}^{2}\right).$ (6)
Algorithm 3 DiverseModel Semi-supervised Learning Pseudocode
  Initialization:
               Randomly initialise 3 models, PSPNet $Q^{1}$, UNet $Q^{2}$, SegNet $Q^{3}$
  INPUT: Labelled Training Dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$
               Unlabelled Training Dataset $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$
  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
        $\mathcal{L}_{sup}=loss(Q^{1}(x_{k}),y_{k})$ + $loss(Q^{2}(x_{k}),y_{k})$
                  + $loss(Q^{3}(x_{k}),y_{k})$ based on (1)
        $\hat{r}_{k}^{1}\leftarrow argmax(Q^{3}(u_{k}))$
        $\hat{r}_{k}^{2}\leftarrow argmax(Q^{3}(u_{k}))$
        $\hat{r}_{k}^{3}\leftarrow argmax(Q^{3}(u_{k}))$
        $\mathcal{L}_{unsup}=loss(Q^{1}(u_{k}),\hat{r}_{k}^{2})$ + $loss(Q^{1}(u_{k}),\hat{r}_{k}^{3})$
                  + $loss(Q^{2}(u_{k}),\hat{r}_{k}^{1})$ + $loss(Q^{2}(u_{k}),\hat{r}_{k}^{3})$
                  + $loss(Q^{3}(u_{k}),\hat{r}_{k}^{1})$ + $loss(Q^{3}(u_{k}),\hat{r}_{k}^{2})$
        $\mathcal{L}\leftarrow\mathcal{L}_{sup}+\lambda\mathcal{L}_{unsup}$
        Update model to minimize $\mathcal{L}$
  OUTPUT: Trained Model $Q^{1}$, $Q^{2}$, $Q^{3}$
The cross-pseudo supervision among three networks creates six losses in the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the average of the six individual losses, as shown below
$\begin{array}[]{r}\mathcal{L}_{unsup}=\frac{1}{6}\left(\mathcal{L}_{unsup}^{12}+\mathcal{L}_{unsup}^{13}+\mathcal{L}_{unsup}^{21}+\mathcal{L}_{unsup}^{23}+\mathcal{L}_{unsup}^{31}+\mathcal{L}_{unsup}^{32}\right).\end{array}$ (7)
The total loss $\mathcal{L}$ is the linear addition of $\mathcal{L}_{sup}$ and $\mathcal{L}_{unsup}$, which is previously given in equation (4)
5 Dataset Description
We employed diverse remote sensing datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset (Rottensteiner et al., 2012), (2) the DFC2020 dataset (Robinson et al., 2021), (3) the RoadNet dataset (Liu et al., 2018), and (4) the Massachusetts Buildings dataset (Mnih, 2013). In the sequel, we share the details of each dataset utilised in this paper.
1) ISPRS Potsdam Semantic Labeling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS). The ground sampling distance of the two modalities of the true orthophoto (TOP) and the DSM is 5 cm. This dataset was annotated manually into six land cover classes, which are impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background. This dataset provides 38 patches (all of the size $6000\times 6000$ pixels), which contain infrared (IR), red, green and blue bands orthorectified optical images with corresponding digital surface models (DSM). For computation purposes, we partitioned all these data tiles into $512\times 512$ patches, resulting in 3456, 201, and 1815 samples for training, validation, and test sets, respectively. We randomly select a quarter of the training data as labelled data and use the remaining three quarters as unlabelled data for semi-supervised learning.
2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset which is based on the SEN12MS dataset (Schmitt et al., 2019). It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps across seven areas namely Mexico City, Kippa-Ring, Khabarovsk, Black Forest, Mumbai, Cape Town, Bandar Anzali between 2016 and 2017. The size of all patches is $256\times 256$ pixels. There are 6112, 986 and 5127 images for training, validation and test sets, respectively. In this paper, we utilize only one-fifth of the labelled data for training, while the remaining four-fifths are employed as unlabeled data. The fine-grained International Geosphere-Biosphere Program (IGBP) classification scheme in SEN12MS was aggregated into ten coarser-grained classes.
3) RoadNet is a benchmark dataset for road detection with 0.21-m spatial resolution per pixel covering 21 typical urban regionsThe annotation data includes road surfaces, road edges, and road centerlines for each image. We only used road surfaces for the segmentation task. the number of samples for training, validation and testing is 410, 45 and 387, respectively. In the semi-supervised learning approaches, only a quarter of the annotated training data is used, while the remaining three-quarters of the training data are employed as unlabeled data for training purposes.
4) Massachusetts Buildings Dataset predominantly encompasses urban and suburban regions, encompassing structures of varying scales. This includes diverse buildings such as individual houses and garages, all of which are included in the labels. The dataset consists of 151 aerial images capturing the Boston area covering roughly 340 square kilometres, with a resolution of $1\leavevmode\nobreak\ m^{2}$ per pixel. The dataset was divided into subsets as follows: 137 images for training, 10 images for testing, and 4 images for validation. Each subset consists of images sized $1500\times 1500$ pixels. A quarter of the annotated training images and the remaining three-quarters of the unlabelled training data are used in semi-supervised learning.
6 Experimental analysis
6.1 Implementation Details
We implemented our methods by using the PyTorch framework. Following (Chen et al., 2018), we used a mini-batch SGD optimizer and adopted a poly learning rate policy, where the current learning rate equals the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$. The initial learning rate and power are set to 0.01 and 0.9, respectively. All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  (McIntosh-Smith, 2014).
We implemented various classic semi-supervised learning methods (MT, CCT, GCT and CPS) on remote sensing datasets. All these methods use DeepLabv3+ with ResNet-50 backbone pre-trained on ImageNet as the semantic segmentation network. To be fair the proposed DiverseHead is also based on DeepLabv3+ with the pretrained ResNet-50 on ImageNet.
We comprehensively analyzed all methods by quantifying performance via class-related measures, including overall accuracy (OA), user’s accuracy (UA), producer’s accuracy (PA), mean intersection over union (mIoU), and $F_{1}$-score. Expressions of all five performance metrics are given as follows: $\mathrm{OA}=\frac{TP+TN}{TP+TN+FP+FN}$, $\mathrm{UA}=\frac{TP}{TP+FP}$, $\mathrm{PA}=\frac{TP}{TP+FN}$, $\mathrm{mIoU}=\frac{|\mathrm{TP}|}{|\mathrm{TP}+\mathrm{FN}+\mathrm{FP}|}$, $F_{1}=\frac{2\cdot PA\cdot UA}{PA+UA}$, where TP, TN, FP, and FN refer to the numbers of pixels that are true positives, true negatives, false positives, and false negatives for each class, respectively.
6.2 Quantitative Results and Analysis
We evaluated the performance of the proposed approaches and some existing semi-supervised frameworks using the five performance metrics above. Experiments are conducted on the four previously mentioned datasets, namely Potsdam, DFC2020, RoadNet, and Massachusetts Building. The average results are presented in Table 1. The reason we first share average results across datasets is to provide a global performance for all utilised techniques, and the detailed performance comparison for all methods on each dataset will be discussed later. From the average results in Table 1, the proposed DiverseHead (DF) achieves the best overall performance in which there are 4 of 5 metrics reach the best compared to other methods, which are marked as red. DiverseModel shows the second-best performance among all methods in Table 1. In particular, for metrics of UA, the proposed DiverseModel exhibits an improvement of over 3.6% compared to another cross-model-based approach, CPS. Although the performance of DiverseHead (DT) is marginally lower than that of DiverseHead (DF) and DiverseModel, the three methods exhibit similar performance and surpass other compared methods. With these used remote sensing datasets, MT demonstrates notably inferior average performance across various metrics. CCT, GCT, and CPS exhibit comparable performance, yet the performance superiority of CPS is notably evident in its PA metric.
Table 1: Average performance comparison with the state-of-the-art methods on four datasets. DT and DF indicate dropout and dynamic freezing, respectively.
Models OA UA PA mIoU $F_{1}$
MT (Tarvainen and Valpola, 2017) 86.33% 74.90% 81.47% 65.97% 77.95%
CCT (Ouali et al., 2020) 87.10% 76.61% 82.29% 67.71% 79.19%
GCT (Ke et al., 2020) 87.43% 75.53% 82.04% 67.19% 78.63%
CPS (Chen et al., 2021) 88.03% 75.90% 85.17% 68.10% 80.13%
DiverseModel (Ma et al., 2023) 88.77% 79.54% 85.18% 70.92% 82.02%
DiverseHead (DT) 88.51% 78.40% 84.84% 70.07% 81.29%
DiverseHead (DF) 88.97% 78.99% 85.82% 71.14% 82.08%
It is important to highlight that the DiverseHead is a significantly lightweight semi-supervised learning architecture aimed at segmentation tasks within remote sensing imagery. Table 2 presents the required number of parameters during training for each semi-supervised learning architecture. Apart from DiverseHead, all other semi-supervised learning approaches typically involve parameters exceeding three hundred megabytes during training. Aside from DiverseModel, all other methods are based on the same segmentation network namely DeepLabv3+. DiverseModel is the largest architecture although it provides competitive performance compared to other cross-model-based methods namely MT, CCT, GCT, and CPS as shown in Table 1. It is evident that the architectures using multiple segmentation networks during training typically need higher memory demands. In contrast, the proposed DiverseHead is remarkably lightweight as it only employs a single model with multiple heads which consists of only a few small convolutional layers, eliminating the need for multiple models during training. Thus, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table 2), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network. Although just a a few parameters are required, the DiverseHead proposed method outperforms this state-of-the-art with at least 1% accuracy and 3% mIoU whilst obtaining similar performance to the companion method of DiverseModel.
Table 2: The required parameter size in each semi-supervised learning approach. Base means supervised segmentation network DeepLabv3+. DT and DF present dropout and dynamic freezing, respectively.
CPS (Chen et al., 2021) MT (Tarvainen and Valpola, 2017) CCT (Ouali et al., 2020) GCT (Ke et al., 2020)
Size (MB) 303.378 303.378 337.655 335.049
DiverseModel (Ma et al., 2023) DiverseHead(DT) DiverseHead(DF) Base
Size (MB) 936.262 175.806 175.806 151.689
Specifically, we present the results of each semi-supervised learning method for the 4 utilised datasets in Table 3. The performance of the DiverseNet family, namely DiverseHead and DiverseModel, is noticeably superior to that of the other listed semi-supervised learning methods across all evaluated datasets, as indicated by the results. In particular, for the Potsdam and Massachusetts Building datasets, the proposed DiverseHead attains the highest performance across 4 out of 5 segmentation metrics whilst DiverseModel emerges as the second-best method based on its results. However, in the case of the RoadNet and DFC2020 datasets, DiverseModel reaches the best for most performance metrics, securing the top position whilst DiverseHead closely follows as the second-best method, despite their similar performance. The performance of DiverseHead (DT) consistently falls between that of DiverseModel and DiverseHead (DF) for most used datasets apart from DFC2020. This dataset provides very low-resolution labels, distinguishing it from other datasets. This could be one of the reasons why DiverseHead (DT) outperforms DiverseHead (DF) in DFC2020.
Table 3: Performance comparison with the state of the arts on four datasets. DT and DF present Dropout and dynamic freezing, respectively.
Model OA UA PA mIoU $F_{1}$
Potsdam MT(Tarvainen and Valpola, 2017) 81.98% 73.66% 78.39% 63.07% 75.95%
CCT(Ouali et al., 2020) 82.66% 74.64% 77.62% 64.16% 76.10%
GCT(Ke et al., 2020) 83.99% 75.65% 80.81% 65.80% 78.14%
CPS(Chen et al., 2021) 85.00% 75.76% 82.94% 66.69% 79.19%
DiverseModel(Ma et al., 2023) 85.76% 76.75% 83.45% 67.85% 79.96%
DiverseHead (DT) 84.66% 77.04% 80.78% 67.12% 78.87%
DiverseHead (DF) 85.98% 79.15% 82.87% 69.63% 80.97%
DFC2020 MT(Tarvainen and Valpola, 2017) 78.64% 59.59% 73.57% 50.44% 65.85%
CCT(Ouali et al., 2020) 79.71% 59.87% 76.40% 51.04% 67.13%
GCT(Ke et al., 2020) 80.84% 61.47% 71.43% 52.17% 66.07%
CPS(Chen et al., 2021) 81.49% 61.74% 79.46% 53.20% 69.49%
DiverseModel(Ma et al., 2023) 81.87% 62.20% 80.69% 53.71% 70.25%
DiverseHead (DT) 82.02% 62.13% 80.61% 53.81% 70.18%
DiverseHead (DF) 81.78% 61.81% 80.21% 53.46% 69.82%
RoadNet MT(Tarvainen and Valpola, 2017) 94.53% 86.77% 87.94% 78.81% 87.35%
CCT(Ouali et al., 2020) 95.45% 89.15% 89.91% 81.97% 89.53%
GCT(Ke et al., 2020) 95.20% 86.27% 90.93% 80.36% 88.54%
CPS(Chen et al., 2021) 95.66% 88.84% 90.95% 82.46% 89.88%
DiverseModel(Ma et al., 2023) 96.84% 93.12% 92.58% 87.11% 92.84%
DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
Massachusetts MT(Tarvainen and Valpola, 2017) 90.16% 79.58% 85.97% 71.55% 82.65%
CCT(Ouali et al., 2020) 90.59% 82.77% 85.23% 73.67% 83.98%
GCT(Ke et al., 2020) 89.67% 78.75% 85.01% 70.42% 81.76%
CPS(Chen et al., 2021) 89.95% 77.26% 87.33% 70.07% 81.98%
DiverseModel(Ma et al., 2023) 90.62% 86.07% 84.01% 75.00% 85.03%
DiverseHead (DT) 91.21% 82.73% 86.98% 74.67% 84.80%
DiverseHead (DF) 91.43% 83.00% 87.45% 75.16% 85.17%
In order to prove the efficiency of the multi-head framework, we considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS). The segmentation framework consists of only a single head. During each training iteration, the current model’s predictions for unlabeled data are used to generate pseudo labels through an $argmax$ operation. These generated pseudo labels are then used to calculate the unsupervised loss. Also, we provide the performance of the segmentation network namely DeepLabv3+, called Base in Table 4, only trained by labelled data (part of each dataset) without using unlabelled data. The performance of the Base and SHS, along with the proposed DiverseHead, are presented in Table 4 to provide experimental evidence of the improvement and efficiency of the use of multiple heads in the DiverseHead architecture. While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the Base model, especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads. Specifically, the mIoU of DiverseHead (DF) is 6.91% higher than that of Base and 5.12% higher than that of SHS for Potsdam. While the performance of DiverseHead (DT) is slightly lower than that of DiverseHead (DF), its performance is still much better than that of SHS and Base. Similarly, DiverseHead demonstrates superior performance on the RoadNet dataset compared to SHS and Base, particularly evident in the mIoU metric, where it exhibits a significant improvement of 6.22% over Base.
Table 4: Performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present dropout and dynamic freezing, respectively.
Model OA UA PA mIoU $F_{1}$
Potsdam Base 81.64% 73.86% 76.05% 62.72% 74.94%
SHS 83.43% 74.57% 79.93% 64.51% 77.16%
DiverseHead (DT) 84.66% 77.04% 80.78% 67.12% 78.87%
DiverseHead (DF) 85.98% 79.15% 82.87% 69.63% 80.97%
RoadNet Base 94.86% 88.22% 88.31% 80.11% 88.26%
SHS 95.32% 87.08% 90.82% 80.95% 88.91%
DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
Table 5: Ablation study on the number of heads for DiverseHead with dynamic freezing
                 # of heads OA UA PA mIoU $F_{1}$
Potsdam                       8 85.92% 77.86% 83.86% 68.91% 80.75%
                      10 85.98% 79.15% 82.87% 69.63% 80.97%
                      12 85.70% 78.54% 82.55% 68.96% 80.50%
RoadNet                       8 96.51% 91.36% 92.55% 85.67% 91.95%
                      10 96.68% 91.99% 92.73% 86.33% 92.36%
                      12 96.66% 92.12% 92.58% 86.31% 92.35%
Average                       8 91.22% 84.61% 88.20% 77.29% 86.35%
                      10 91.33% 85.57% 87.80% 77.98% 86.66%
                      12 91.18% 85.33% 87.57% 77.64% 86.42%
An ablation study was conducted on the Potsdam and RoadNet datasets to explore the impact of the number of heads for the proposed DiverseHead method with dynamic freezing perturbation. The results in Table 5 indicate that using 10 heads yields the optimal choice across most metrics in terms of average performance though the performance difference between the versions with different numbers of heads is small on both datasets. Thus, we chose the 10-head version of DiverseHead for all experiments in this paper.
Table 6: Performance comparison of the DiverseModel with its constituent networks on Potsdam dataset
Model OA UA PA mIoU $F_{1}$
UNet 83.28% 74.18% 78.88% 64.51% 76.46%
SegNet 82.37% 73.53% 77.46% 63.24% 75.45%
PSPNet 81.96% 74.23% 76.99% 63.08% 75.58%
DiverseModel 85.76% 76.75% 83.45% 67.85% 79.96%
Since the DiverseModel technique uses three different segmentation networks, we also evaluated the performance of each member network within the DiverseModel on the Potsdam dataset in Table 6. The component models UNet, SegNet, and PSPNet use unlabeled data through the way of SHS. DiverseModel demonstrates a significant improvement across all performance metrics in the segmentation task compared to each network. In particular, the metric of PA experiences a notable improvement of 4.57% compared to the best-performing individual component model. This phenomenon can be attributed to the enhancement of pseudo-label diversity through cross-model supervision, resulting in a significant improvement in PA (recall) during the test phase. The findings suggest that the cross-supervision of different networks has the potential to achieve superior performance compared to the best-performing individual component.
Potsdam DFC2020 RoadNet Massachusetts
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IoU = 39.38% IoU = 70.08% IoU = 79.30% IoU = 73.37%
CCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IoU = 51.19% IoU = 78.21% IoU = 86.24% IoU = 75.97%
GCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IoU = 40.35% IoU = 76.66% IoU = 80.96 IoU = 74.20%
Potsdam DFC2020 RoadNet Massachusetts
CPS Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 50.21% IoU = 79.39% IoU = 84.92% IoU = 74.16%
DiverseModel Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 58.76% IoU = 80.28% IoU = 94.67% IoU = 76.84%
DiverseHead (Dropout) Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 57.46% IoU = 80.12% IoU = 93.01% IoU = 78.93%
DiverseHead (Dynamic Freezing) Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 58.25% IoU = 80.31% IoU = 90.35% IoU = 78.17%
Figure 7: Visual results of each method on four datasets. The IoU score of prediction for each method is shown in its sub-caption.
6.3 Qualitative Results and Analysis
The visible results are presented in Figures 7. For each dataset, we randomly selected one case of the RGB image, its ground truth, and predictions for all the methods. Also, the IoU scores are calculated for all predictions with the ground truth for each case, which are shown in the sub-caption of each prediction. Based on the IoU values of visual predictions, the highest score is obtained by either DiverseModel or DiverseHead (DT&DF). Especially for the RoadNet dataset, both the DiverseModel and DiverseHead families achieve an IoU of over 90% surpassing other methods by at least 3.5%. Visually, the segmentation maps of DiverseModel and DiverseHead display better continuity and show better similarity to the ground truth than other methods.
7 Conclusion
In this paper, we proposed a lightweight semi-supervised learning approach based on a multi-head structure called DiverseHead. Based on the multi-head structure we provide two perturbation methods, namely dynamic freezing and dropout. Taking inspiration from the theory of bagging, a voting mechanism is proposed to generate beneficial pseudo-labels in the training stage. This simple and lightweight semi-supervised learning framework shows competitive performance for the segmentation of remote sensing imagery. Furthermore, we conducted further analysis and additional evaluation on the previously proposed multi-network-based semi-supervised learning method known as DiverseModel. Based on the results obtained from the aforementioned four remote sensing datasets, DiverseHead and DiverseModel demonstrate comparable performance while significantly outperforming various widely used semi-supervised learning frameworks. It is important to highlight that the training structure of DiverseHead is significantly lighter than that of DiverseModel and other state-of-the-art methods, while still achieving competitive performance. This implies that DiverseHead is a better option for those who lack high-memory computation resources.